OpenStudy (narfthepanda):

find the quotient -7/ 14 over 3 im stumped would appreciate the help

OpenStudy (anonymous):

So, is this what you want? \[\frac{-7/14}{3}\]

OpenStudy (narfthepanda):

yes

OpenStudy (anonymous):

First, let's simplify a bit... what's a simpler form for -7/14 ?

OpenStudy (narfthepanda):

ummmm something having to do with 2?

OpenStudy (narfthepanda):

im sorry im not good at math at all

OpenStudy (anonymous):

I want you to do something like this (here's a different example): 6/12 = 1/2 You think: 6 goes into 6 once; 6 goes into 12 twice. So ... what's the simplest form for -7/14?

OpenStudy (narfthepanda):

it would also be one half no?

OpenStudy (anonymous):

Don't apologize! It's so awesome that you're here trying to understand things!

OpenStudy (narfthepanda):

1/2

OpenStudy (anonymous):

Close --- you're just missing the minus sign in front: -7/14 = -1/2

OpenStudy (narfthepanda):

ooooooo ok

OpenStudy (anonymous):

So far, this is what we have: \[\frac{-7/14}{3} = \frac{-1/2}{3}\] Now, we need to remember: dividing by 3 is the same as multiplying by 1/3 . Do you believe that?

OpenStudy (narfthepanda):

yes

OpenStudy (anonymous):

So, \[\frac{-1/2}{3} = -\frac 12\cdot \frac{1}{3}\] Can you take it from here?

OpenStudy (narfthepanda):

yes thank you sooooooo much

OpenStudy (anonymous):

You're welcome! Happy Friday!

OpenStudy (narfthepanda):

you too

OpenStudy (anonymous):

Thanks!